Circular No. 5776

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 S. D. Van Dyk and K. W. Weiler, Naval Research Laboratory; R. A.

Sramek and M. P. Rupen, National Radio Astronomy Observatory; and N. Panagia, Space Telescope Science Institute, report: “New observations of SN 1993J made on Apr. 22.47 UT with the Very Large Array in ‘B’ configuration show that the source currently has a flux density at 1.3 cm (22.5 GHz) of 39.94 ± 0.53 mJy; the radio light curve at this frequency appears to be flattening as expected from preliminary model predictions (IAUC 5762). A slow decline in the flux density at 1.3 cm should commence shortly. The supernova also has flux densities of 37.35 ± 0.34 mJy at 2.0 cm (14.9 GHz), 11.77 ± 0.07 mJy at 3.6 cm (8.4 GHz), and 1.70 ± 0.07 mJy at 6 cm (4.86 GHz). This flux density evolution is consistent with previous model estimates (IAUC 5762) to within the predictive capability of the model from the few data points available at that time. Revised estimates will be issued shortly. Radio monitoring is continuing on a nearly daily basis at these and other wavelengths in order to follow the supernova’s rapidly changing behavior.”

B. Jannuzi, Institute for Advanced Study; G. Schmidt, Steward Observatory; R. Elston, Cerro Tololo Interamerican Observatory; and P. Smith, Steward Observatory, report: “Optical polarization was detected from SN 1993J on Apr. 26.12 UT using the CCD spectropolarimeter of Steward Observatory on the Kitt Peak Mayall 4-m telescope. Over the wavelength interval 400-735 nm, the polarization averages 1.04 percent at a position angle of 29^{o}, but considerable structure is present through the strong emission lines. Dilution of the continuum polarization is apparent at Hα, but a 10^{o} rotation in position angle through the middle of the remarkable doublepeaked profile implies that the line flux is intrinsically polarized. Polarization is a new phenomenon in SN 1993J, as P. Smith obtained a V -band value of 0.19 ± 0.11 percent at 178^{o} ± 17^{o} on Apr. 7.22 UT using the Two-Holer polarimeter on the Steward Observatory 1-m telescope. Observations of nearby stars confirm that interstellar polarization in the direction of the supernova is less than 0.1 percent.”

PERIODIC COMET SCHWASSMANN-WACHMANN 1 CCD V total magnitudes from H. Mikuž, Ljubljana, Slovenia (0.20-m

Baker-Schmidt camera): Apr. 8.82 UT, 13.3; 22.83, 12.6; 23.84, 12.8.

1993 April 28 Daniel W. E. Green